# Aethorix v1.0: An Integrated Scientific AI Agent for Scalable Inorganic Materials Innovation and Industrial Implementation

### v3

@@ 6 Structure Generation
All atomistic structural modeling is managed by the Atomic Simulation Environment (ASE) [larsen2017atomic]. The source of initial structures is contingent upon the design objective: for targeted exploration of known crystal systems, initial configurations are sourced from the Materials Project[jain2013commentary] database with rational compositional and configurational modifications, whereas for dede novonovo design, the goal is accomplished using a symmetry-aware, diffusion-based generative model. The implementation is adapted from MatterGen[zeni2025generative], which is architected for zero-shot generation of novel and thermodynamically stable inorganic materials. In this scheme, periodic crystal materials are represented by $M=(A,X,L)$, where $A=(a^{1},a^{2},...,a^{n})^{T}\in\mathbb{A}^{n}$ represents the atom species inside the unit cell, $X=(x^{1},x^{2},...,x^{n})\in[0,1)^{3\times n}$ denotes the fractional coordinates of atoms in the unit cell, and $L=(l^{1},l^{2},l^{3})\in\mathbb{R}^{3\times 3}$ is the lattice with three vector components. During the forward diffusion (noising) process, AA, $X$, and $L$ are independently corrupted toward a physically meaningful prior distribution as follows:
@@ 7 Structure Optimization
where $\alpha_{E}$, $\alpha_{F}$, and $\alpha_{S}$ are the weight coefficients for total energy, atomic force, and virial stress, respectively, NN denotes the number of training samples, $X_{i}$ and $\hat{X}_{i}$ ($X=E,\mathbf{F},\sigma$) are the true and predicted targets for structure $i$, respectively. Optimized hyperparameters are often determined using Bayesian optimization methods.[martinez2018practical] MLIP ensembles are constructed by training multiple models on identical datasets but with different parameter initializations. The variance in predictions across the ensemble provides a rigorous, $a$ prioripriori metric for assessing model confidence and identifying extrapolation beyond the trained domain[musil2019fast]. Active learning is implemented to ensure transferability of specialized models across the desired chemical space.[podryabinkin2017active, podryabinkin2019accelerating, zhang2019active] The training dataset is iteratively enriched by incorporating new structures that are generated from the conditioned atom types and have the highest epistemic uncertainty.
@@ 8 Physics-Informed Prediction
Many emergent macroscale properties of materials cannot be simply derived from a single DFT geometric optimization calculation, but instead require multiple DFT calculations integrated with physics-informed analysis. Hence, the computational cost can grow exponentially with increasing system size and project scope. In contrast, MLIPs can unlock the unprecedented efficiency in predictive screening while preserving near-abab initioinitio fidelity. To enable higher-level property predictions, our current version provides direct interfaces to Pymatgen[ong2013python], Phonopy, Phono3py[togo2023first, togo2023implementation], and DeePMD-kit.[zeng2023deepmd] Figure 2 contrasts the computational cost scaling between DFT and MLIPs across a range of system sizes for various simulation tasks. These performance benchmarks are averaged from the estimated calculation convergence time observed across a series of Mg-Si-O-H chemical systems. The computational cost of obtaining emergent properties via MLIP inference is 103 $\sim$ 105 times lower than equivalent DFT calculations. It should be noted that the exact magnitude of this advantage can vary with the chemical system, the resolution of the computational methods employed ($e.g.,$ the number of intermediate images in nudged elastic band calculations), and the simulation timescale ($e.g.,$ for molecular dynamics trajectories), yet it consistently reflects a significant reduction in computational expense for MLIP. Crucially, this cost reduction renders many prohibitively expensive calculations ($e.g.,$ phonon vibrations) feasible for high-throughput screening.
@@ 9.2 Phase Stabilization through Atomistic Engineering
where NN is the number of atoms, $d$ is the dimensionality of diffusion, $\mathbf{r}_{i}$ is the position of atom $i$ at $t$, and the angular brackets denote an ensemble or time average. The self-diffusivity was found to vary by 1–2 orders of magnitude across all thermodynamically most stable polymorphs. Those with low ionic mobility ($e.g.,$ ID 01) were hence identified undesirable and their production was targeted for suppression during synthesis. Figure 4g and S1c show the phase stabilization effect due to presence of low concentration ($\sim$10 mol%) of dopant ions in clinker crystal lattices at 1730 K. The results revealed varying dopant site occupancy preferences across polymorphs, which deviate from empirical trends.[zhang2011development, young1996highly, schwiete1968existenzbereiche] For instance, Na preferentially substitutes the Ca site in the C2S ID10 lattice, whereas it favors the Si site in the C3S ID25 lattice. Most of these substitutions resulted in a reduced free energy at 1730 K, Notable exceptions, such as S or P substituting for Ca, are likely due to the significant charge mismatch introduced by the dopant, which destabilizes the lattice. These are supported by experimental observations.[ludwig2015research] Crucially, K doping provided the strongest stabilization of all dopants for the desired polymorphs, while having a negligible effect on the undesired ones. It was further verified that incorporation of K does not impede ion mobility within the desired phases. Therefore, it was proposed that industrial clinker production strategically maintain K impurity levels at 5–10 mol% to favor the formation of hydration-reactive polymorphs, a direct route to improving the performance of concrete from early ages to long-term service.

### v4

@@ 6 Structure Generation
All atomistic structural modeling is managed by the Atomic Simulation Environment (ASE) [larsen2017atomic]. The source of initial structures is contingent upon the design objective: for targeted exploration of known crystal systems, initial configurations are sourced from the Materials Project[jain2013commentary] database with rational compositional and configurational modifications, whereas for dede novonovo design, the goal is accomplished using a symmetry-aware, diffusion-based generative model. The implementation is adapted from MatterGen[zeni2025generative], which is architected for zero-shot generation of novel and thermodynamically stable inorganic materials. In this scheme, periodic crystal materials are represented by $M=(A,X,L)$, where $A=(a^{1},a^{2},...,a^{n})^{T}\in\mathbb{A}^{n}$ represents the atom species inside the unit cell, $X=(x^{1},x^{2},...,x^{n})\in[0,1)^{3\times n}$ denotes the fractional coordinates of atoms in the unit cell, and $L=(l^{1},l^{2},l^{3})\in\mathbb{R}^{3\times 3}$ is the lattice with three vector components. During the forward diffusion (noising) process, $A$, $X$, and $L$ are independently corrupted toward a physically meaningful prior distribution as follows:
@@ 7 Structure Optimization
where $\alpha_{E}$, $\alpha_{F}$, and $\alpha_{S}$ are the weight coefficients for total energy, atomic force, and virial stress, respectively, $N$ denotes the number of training samples, $X_{i}$ and $\hat{X}_{i}$ ($X=E,\mathbf{F},\sigma$) are the true and predicted targets for structure $i$, respectively. Optimized hyperparameters are often determined using Bayesian optimization methods.[martinez2018practical] MLIP ensembles are constructed by training multiple models on identical datasets but with different parameter initializations. The variance in predictions across the ensemble provides a rigorous, $a$ prioripriori metric for assessing model confidence and identifying extrapolation beyond the trained domain[musil2019fast]. Active learning is implemented to ensure transferability of specialized models across the desired chemical space.[podryabinkin2017active, podryabinkin2019accelerating, zhang2019active] The training dataset is iteratively enriched by incorporating new structures that are generated from the conditioned atom types and have the highest epistemic uncertainty.
@@ 8 Physics-Informed Prediction
Many emergent macroscale properties of materials cannot be simply derived from a single DFT geometric optimization calculation, but instead require multiple DFT calculations integrated with physics-informed analysis. Hence, the computational cost can grow exponentially with increasing system size and project scope. In contrast, MLIPs can unlock the unprecedented efficiency in predictive screening while preserving near-$ab$ initioinitio fidelity. To enable higher-level property predictions, our current version provides direct interfaces to Pymatgen[ong2013python], Phonopy, Phono3py[togo2023first, togo2023implementation], and DeePMD-kit.[zeng2023deepmd] Figure 2 contrasts the computational cost scaling between DFT and MLIPs across a range of system sizes for various simulation tasks. These performance benchmarks are averaged from the estimated calculation convergence time observed across a series of Mg-Si-O-H chemical systems. The computational cost of obtaining emergent properties via MLIP inference is 103 $\sim$ 105 times lower than equivalent DFT calculations. It should be noted that the exact magnitude of this advantage can vary with the chemical system, the resolution of the computational methods employed ($e.g.,$ the number of intermediate images in nudged elastic band calculations), and the simulation timescale ($e.g.,$ for molecular dynamics trajectories), yet it consistently reflects a significant reduction in computational expense for MLIP. Crucially, this cost reduction renders many prohibitively expensive calculations ($e.g.,$ phonon vibrations) feasible for high-throughput screening.
@@ 9.2 Phase Stabilization through Atomistic Engineering
where $N$ is the number of atoms, $d$ is the dimensionality of diffusion, $\mathbf{r}_{i}$ is the position of atom $i$ at $t$, and the angular brackets denote an ensemble or time average. The self-diffusivity was found to vary by 1–2 orders of magnitude across all thermodynamically most stable polymorphs. Those with low ionic mobility ($e.g.,$ ID 01) were hence identified undesirable and their production was targeted for suppression during synthesis. Figure 4g and S1c show the phase stabilization effect due to presence of low concentration ($\sim$10 mol%) of dopant ions in clinker crystal lattices at 1730 K. The results revealed varying dopant site occupancy preferences across polymorphs, which deviate from empirical trends.[zhang2011development, young1996highly, schwiete1968existenzbereiche] For instance, Na preferentially substitutes the Ca site in the C2S ID10 lattice, whereas it favors the Si site in the C3S ID25 lattice. Most of these substitutions resulted in a reduced free energy at 1730 K, Notable exceptions, such as S or P substituting for Ca, are likely due to the significant charge mismatch introduced by the dopant, which destabilizes the lattice. These are supported by experimental observations.[ludwig2015research] Crucially, K doping provided the strongest stabilization of all dopants for the desired polymorphs, while having a negligible effect on the undesired ones. It was further verified that incorporation of K does not impede ion mobility within the desired phases. Therefore, it was proposed that industrial clinker production strategically maintain K impurity levels at 5–10 mol% to favor the formation of hydration-reactive polymorphs, a direct route to improving the performance of concrete from early ages to long-term service.